OpenStudy (anonymous):

i just want to make sure I am correct on this . Here is my math problem on reducing fractions 20/60 = 10/10 = 2/6= 1/3 is this correct?

OpenStudy (anonymous):

20/60 =2/6=1/3 is doesn't equal to 10/10

OpenStudy (anonymous):

ok then what should it be 20/20 = 1/3

OpenStudy (anonymous):

20/20=1/3?

OpenStudy (anonymous):

20/20=1/1 =1

OpenStudy (anonymous):

so the answer is 1 correct. so 24/32=8/8=/34 correct

OpenStudy (anonymous):

that 3/4

myininaya (myininaya):

how are you getting everything equals one?

OpenStudy (anonymous):

\[ 32= (2*2*2*2*2)\] \[24=2*2*2*3\]

myininaya (myininaya):

anything you have the same number over the same number it is one

OpenStudy (anonymous):

24/32 = 2*2*2*3/(2*2*2*2*2)= 3/(2*2)=3/4

OpenStudy (anonymous):

ok so my answer to the problem 24/34 is 3/4 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you can't write 24/34=8/8

myininaya (myininaya):

oh i know what he meant now hes dividing top and bottom by 8

OpenStudy (anonymous):

ok. but the way my teacher wrote the example was 21/28 = 7/7= 3/4

myininaya (myininaya):

don't do it. its incorrect. and your teacher did she/he is incorrect

OpenStudy (anonymous):

21/28 * 7/7 maybe?

OpenStudy (anonymous):

maybe it was a type incorrect. let me try on more. 50/75 =2/3

myininaya (myininaya):

correct!

OpenStudy (anonymous):

so 31/93 = 1/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

how about adding fractions? can you help me with that

OpenStudy (anonymous):

6/11 + 3/22 = would that be 9/33 = 3/11

OpenStudy (anonymous):

I would assume the teacher meant to write divided by 7/7 you can divide any number by one and it does not change its value.

OpenStudy (radar):

6/11 +3/22 is not 9/33 6/11 X 2/2 =12/22 now you can add 12/22 + 3/22= 15/22

OpenStudy (radar):

Although if it was 9/33 it does equal 3/11